# When Fairness Meets Privacy: Exploring Privacy Threats in Fair Binary Classifiers through Membership Inference Attacks

### v2

@@ 2 Related Work
Algorithmic fairness methods are categorized into pre-processing, in-processing, and post-processing approaches based on their processing stage. We focus on in-processing approaches as they modify model training procedures directly and may pose model privacy threats. Among them, some introduce fair constraints and formulate the issues as optimization problems Zemel et al. (2013); Manisha and Gujar (2020); Tang et al. (2023); Truong et al. (2023); Cruz et al. (2023); Jung et al. (2023). Some propose adversarial designs to remove sensitive information among extracted features Zhu et al. (2021); Creager et al. (2019); Park et al. (2021). Some adopt data sampling Roh et al. (2021) or reweighting Chai and Wang (2022) approaches to alleviate the biased predictions. More recently, studies learn fair representations using mixup augmentations or contrastive learning methods Ching-Yao Chuang (2021); Du et al. (2021); Park et al. (2022); Wang et al. (2022); Zhang et al. (2023a); Qi et al. (2022). These methods interpolate inputs or modify features to pursue fair representations.
@@ 4.1 Naive score-based attacks
Moreover, we explore the score changes for different subgroups in Figures 4c and 4d. From the plots, the majority data tend to be more “spread out”, whereas the minority are “more concentrated”. This is because fairness interventions strive to balance prediction performance across subgroups for fair predictions. The results advocate the observed increased loss values for most data points in Figure 1a This observation also aligns with the fairness-utility trade-off, extensively observed in prior studies Zhang et al. (2023b); Pinzón et al. (2022); Zietlow et al. (2022), where achieving fairness often compromises model prediction performance.

### v3

@@ 2 Related Work
Algorithmic fairness methods are categorized into pre-processing, in-processing, and post-processing approaches based on their processing stage. We focus on in-processing approaches as they modify model training procedures directly and may pose model privacy threats. Among them, some introduce fair constraints and formulate the issues as optimization problems Zemel et al. (2013); Manisha and Gujar (2020); Tang et al. (2023); Truong et al. (2023); Cruz et al. (2023); Jung et al. (2023). Some propose adversarial designs to remove sensitive information among extracted features Zhu et al. (2021); Creager et al. (2019); Park et al. (2021). Some adopt data sampling Roh et al. (2021) or reweighting Chai and Wang (2022) approaches to alleviate the biased predictions. More recently, studies learn fair representations using mixup augmentations or contrastive learning methods Ching-Yao Chuang (2021); Du et al. (2021); Park et al. (2022); Wang et al. (2022); Zhang et al. (2023); Qi et al. (2022). These methods interpolate inputs or modify features to pursue fair representations.
@@ 4.1 Naive score-based attacks
Moreover, we explore the score changes for different subgroups in Figures 4c and 4d. From the plots, the majority data tend to be more “spread out”, whereas the minority are “more concentrated”. This is because fairness interventions strive to balance prediction performance across subgroups for fair predictions. The results advocate the observed increased loss values for most data points in Figure 1a This observation also aligns with the fairness-utility trade-off, extensively observed in prior studies Zhang et al. (2023); Pinzón et al. (2022); Zietlow et al. (2022), where achieving fairness often compromises model prediction performance.
@@ 6 Mitigation
7 Discussions
7.1 Influential factors
The proposed FD-MIA exploits the prediction gaps between biased and fair models to achieve enhanced MIA results. As indicated in Figures 8 and 9 of the main paper, the attack results can be influenced by two factors: the adopted fair model and the fairness method used.
In the main experiments, we considered fair models without severe accuracy degradation after fairness interventions. This led to smaller prediction gaps between the fair and biased models, resulting in relatively modest improvements in the attack performance.
Figure 8 shows that FD-MIA adopts fair models with varying fairness performance. We observe that fairer models generally lead to better attack results. This is due to the enlarged prediction gaps between the biased and fair models.
Figure 9 considers different fairness methods, which result in different attack results. This discrepancy indicates that the fairness methods can also influence the attack results. Notably, the adversarial training and constraint approaches lead to more pronounced prediction gaps between biased and fair models, resulting in enhanced attack results. This may be due to the more substantial trade-offs between fairness and model utility inherent to these approaches.
7.2 Ethical Considerations
We reveal a concerning tension between the pursuit of algorithmic fairness and the preservation of individual privacy. While fairness interventions have shown promise in mitigating biased predictions, our results demonstrate that these techniques can inadvertently introduce new vulnerabilities to membership inference attacks, posing significant risks to model privacy.
This trade-off between fairness and privacy raises important ethical questions that warrant careful examination. On one hand, there is a moral imperative to ensure machine learning models do not perpetuate harmful biases and discriminatory practices On the other hand, the privacy of individuals whose data is used to train these models is also an essential ethical consideration.
To explore the trade-offs between fairness and privacy, we measure model fairness using different metrics and assess the model’s privacy performance with MIAs. These practices can be adopted to evaluate deployed models. The Model developers should engage with ethicists, policymakers, and affected communities to ensure that there are mechanisms in place to hold these systems accountable for their fairness and privacy implications.
To mitigate the privacy risks while maintaining fairness, we encourage the adoption of techniques like differential privacy (DP-SGD). By introducing carefully calibrated noise into the training process, differential privacy can help preserve the privacy of individuals whose data is used to train the models without significantly compromising their fairness performance.
Finally, the trade-off between fairness and privacy is a complex issue that requires nuanced ethical deliberation. By incorporating these key considerations, we aim to provide practitioners with a more holistic understanding of the challenges and guide them in making informed decisions that uphold both ethical principles to the greatest extent possible.